OpenStudy (anonymous):

Solve using the elimination method. Show your work. If the system has no solution or an infinite number of solutions, state this. -6x – 6y = -12 -6x + 7y = 14

OpenStudy (anonymous):

-6x – 6y = -12---------1 -6x + 7y = 14 ---------2 subtract 2nd equation from 1st equation -6x – 6y = -12 +6x - 7y =-14 -------------- -13y=-26 y=-26/-13 y=2 now plug in y value into one of the equation -6x+7(2)=14 -6x+14=14 -6x=14-14 x=0 so,x=0 and y=2